You find a zero coupon bond with a par value of $15,000 and 20 years to maturity. If the yield to maturity on this bond is 5.8 percent, what is the dollar price of the bond? Assume semiannual compounding periods.
Answer #1

Price of a zero coupon bond is equal to the discounted value of its face value

Semi-annual required return = 5.8%/2 = 2.9%

Number of semi-annual periods = 20*2 = 40

Dollar price of bond = 15,000*PVF(2.9%, 40 years)

= 15,000*0.3187

= $4,780.5